The no. of 6 digit nos that can be formed from the digits 1,2,3,4,5,6,&7 so that digits do not repeat and the terminal digits are even is?

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear jee

first terminal digit can be filled in 3 ways

second terminal digit can be filled in 2 ways

rest 4 place can be filled in 5*4*3*2 ways

total ways =3*2*5*4*3*2


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

dravesh jaluka
6 years ago
If 1 select and even digit  as terminal digit , I can choose the remaining 5 numbers out of 6 in 6 ways and arrange them in 5! ways i.e 120 ways .
Therefore when even terminal digit I have 6 *120 = 720 options.
Therefore for 3 even terminal digits I have 3 *720 = 2160 options